OpenStudy (anonymous):

if f(x)=7+4x and g(x)=1/2x what is the value of (f/g) (5)

Vocaloid (vocaloid):

(f/g)(5) = f(5)/g(5) f(5) = ?

OpenStudy (ageta):

f5/g5=f5/g5

OpenStudy (ageta):

they are equal

OpenStudy (anonymous):

11/2 160 270 27/10

OpenStudy (anonymous):

are the options

Vocaloid (vocaloid):

f(x) = 7 + 4x f(5) = ?

OpenStudy (anonymous):

so do i put the f(5) where the f(x) is??? im sorry im so confused haha

Vocaloid (vocaloid):

replace "x" with 5

Vocaloid (vocaloid):

f(x) = 7 + 4x f(5) = 7 + 4(5) = ?

OpenStudy (anonymous):

27

Vocaloid (vocaloid):

right, now calculate g(5)

OpenStudy (anonymous):

1/10 so the answer would be 27/10?

Vocaloid (vocaloid):

not quite f(5)/g(5) = ?

OpenStudy (anonymous):

ohhhhh 270 :)

OpenStudy (anonymous):

thank you!!!!!

Vocaloid (vocaloid):

right